2. (20 points) (a). Find the derivative of the complex-valued function f() = (2-21) (3z + 5) at z = 31.
Answer #1

Solution! Given complexed - valued function foz) = (2-21) (3275) Cluim. find derivative at point 2:31 Now using the formula f- lim A230 12142 43(42) 24542 Az - lim i sh Cl2 i +3 (A2) 15) A210 sk . 21 +3 (A2) t5 , a lim 42so Put sz=2-31 cim2 isi i 1 3